The Peace Barber Shop employs four barbers. One barber, who also serves as the manager, is paid a salary of $1,830 per month. The other barbers are paid $1,350 per month. In addition, each barber is paid a commission of $2 per haircut. Other monthly costs are as follows: store rent $830 plus 60 cents per haircut; depreciation on equipment $520; barber supplies 40 cents per haircut; utilities $320; and advertising $250. The price of a haircut is $13.

Instructions: a) Determine the variable cost per haircut and total monthly fixed costs b)Calculate break-even point in units and dollars c) Determine operating income , assuming 1580 haircuts are given in a month.
Answer #1

A) Variable cost Per Hair Cut = $3.00 (Answer)

Barber Commission +Store rent +barber supplies = 2+0.60+0.40 = $3.00

Fixed cost Per Month : $7800 Per Month

Barber salaries + Store Rent +Depreciation +Utilitites +Advertisement

= 1850+(3*1350)+830+520+320+250 = $7,800

B) Break Even Units = Fixed cost / Contribution Margin

contribution Margin = 13-3 = 10

  • Break Even Units = $7,800 /10 =780 Units
  • Break Even Dollars = 780*13 = $10,140

C) Operating income at 1580 Hair Cuts

Operating income = (1580*10)-7800 = $8,000 (Answer)